Instructions: Enter your responses rounded to the nearest billion.

a. Calculate real GDP for 2016 using 2000 prices.

     $  billion

b. By how much did real GDP increase between 2000 and 2016?

     $  billion

c. By how much did nominal GDP increase between 2000 and 2016?

     $  billionRefer to the data below (all values are in billions): Year Nominal GDP GDP Deflator (in Billions of Dollars) (Price Index 2009 100) 2000 2001 2002 2003 2004 2005 2006 2007 2008 2009 2010 2011 2012 2013 2014 2015 2016 9,817 10,128 10,469 10,960 11,685 12,422 13,178 13,808 14,291 13,939 14,524 15,518 16,163 16,768 17,393 18,037 18,569 81.9 83.8 85.0 86.7 89.1 92.0 94.8 97.3 99.2 100.0 101.2 103.3 105.2 106.7 108.8 110.0
Answer #1

a) Real GDP in 2016 using 2000 prices

= 18569*100/81.9

= 22672.77 billion

b) Real GDP in 2000 = 9817*100/81.9 = 11986.56 billion

Real GDP in 2016 = 18569*100/111.4 = 16668.76 billion

Real GDP increased by 16668.76-11986.56 = 4682.2 billion

c) Nominal GDP increased by 18569-9817 = 8752 billion